# Jogging the Memory of Unlearned LLMs Through Targeted Relearning Attacks Content warning: This manuscript contains examples of harmful/hazardous text

### v2

@@ D.4 Unsuccessful Relearning Example
Appendix E Overview of Unlearning Baselines
@@
E.1 Gradient Ascent and Gradient Difference
@@
Gradient difference [16] adds a term to minimize loss on the retain set at the same time. The objective of gradient difference can be written as
@@
E.2 KL Minimization
KL minimization [21] aims to minimize the KL divergence between the current model and reference model $\mathbf{w}_{\rm ref}$ (model at the start of unlearning) on the retain set $D_{r}$ outputs while performing gradient ascent on the unlearn set $D_{u}$. The objective can be written as
@@ E.2 KL Minimization
E.3 Negative Preference Optimization (NPO)
NPO [37] comes from the idea fitting unlearning problem to preference optimization framework [26]. In this case, we penalizes the prompt-response pairs in the forget set and ignore the positive response. Denote the reference model weights as $\mathbf{w}_{\rm ref}$, the objective can be written as
@@ E.3 Negative Preference Optimization (NPO)
E.4 SCRUB
SCRUB [13] utilizes a combination of minimization of KL divergence between the reference model and the current model on $D_{r}$ outputs, maximization of KL divergence on $D_{u}$ outputs, and gradient descent on $D_{r}$. However, instead of combining all three objectives together, this methods alternative between min and max steps. During the min step, the first and last term is combined i.e.
@@ E.4 SCRUB
E.5 RMU
RMU [14] seeks to degrade unlearn data’s representation within the model. To achieve this goal without harming model performance on retain knowledge, a subset of layers (usually 1 layer) is chosen, and the input representation on such layers are aligned towards noise vectors for $D_{u}$, and reference representations obtained from the reference model for $D_{r}$. The objective can be written as
@@ E.5 RMU
E.6 WHP Alternate Labels
In WHP [7], the unlearning is performed through finetuning on alternate labels. The methods of producing alternate labels consists of two parts. Firstly, certain key terms (called “anchor terms”) are identified and replaced with generic terms according to a pre-set dictionary. For other tokens, this method hopes to identify the tokens that has the most positive change from a baseline model to a reinforced model that has been fitted to the unlearn data. Such increase are reversed according to the formula

### v4

@@ D.4 Unsuccessful Relearning Example
Appendix E RMU experiments
Phase MCQ Acc QA Score
Original 0.5177 5.92
RMU 0.3086 1.89
Relearn 0.4812 2.24
Zero init+Relearn 0.4029 4.84
We evaluate on two tasks for RMU: the same answer completion task in Section 4 and the original MCQ in (Li et al., 2024). The later is evaluated by calculating the log likelihood for each answer instead of asking the model to generate text.
As shown by the first three rows of Table 8, relearning is able to significantly increase the MCQ accuracy compared to the unlearned model, even if the relearn set does not contain direct answer to these MCQ questions by construction. On the other hand, we see little improvement over the LLM-as-Judge score for the completion task. We observe that a majority of relearned model output for the completion task is actually simple repetition of the same token, similar to the output generated from the unlearned model. Hence, despite the relearned model achieves higher MCQ accuracy, it doesn’t mean the model perform well on other tasks (e.g. text generation) on the forget set. To tackle this issue, we come up with a slightly different version of the relearning attack for unlearning methods like RMU. Given that we know the layers being corrupted, we imply re-initialize the layer to be 0 everywhere, and then follow the same recipe to finetune on the relearn set. We call this method zero init+relearn and shows its performance on the last row of Table 8. While our method degrades the model knowledge in MCQ, it improves the answer completion score by a lot via removing the corrupted weights.
Appendix F Overview of Unlearning Baselines
@@
F.1 Gradient Ascent and Gradient Difference
@@
Gradient difference (Liu et al., 2022) adds a term to minimize loss on the retain set at the same time. The objective of gradient difference can be written as
@@
F.2 KL Minimization
KL minimization (Maini et al., 2024b) aims to minimize the KL divergence between the current model and reference model $\mathbf{w}_{\rm ref}$ (model at the start of unlearning) on the retain set $D_{r}$ outputs while performing gradient ascent on the unlearn set $D_{u}$. The objective can be written as
@@ F.2 KL Minimization
F.3 Negative Preference Optimization (NPO)
NPO (Zhang et al., 2024) comes from the idea fitting unlearning problem to preference optimization framework (Rafailov et al., 2024). In this case, we penalizes the prompt-response pairs in the forget set and ignore the positive response. Denote the reference model weights as $\mathbf{w}_{\rm ref}$, the objective can be written as
@@ F.3 Negative Preference Optimization (NPO)
F.4 SCRUB
SCRUB (Kurmanji et al., 2024) utilizes a combination of minimization of KL divergence between the reference model and the current model on $D_{r}$ outputs, maximization of KL divergence on $D_{u}$ outputs, and gradient descent on $D_{r}$. However, instead of combining all three objectives together, this methods alternative between min and max steps. During the min step, the first and last term is combined i.e.
@@ F.4 SCRUB
F.5 RMU
RMU (Li et al., 2024) seeks to degrade unlearn data’s representation within the model. To achieve this goal without harming model performance on retain knowledge, a subset of layers (usually 1 layer) is chosen, and the input representation on such layers are aligned towards noise vectors for $D_{u}$, and reference representations obtained from the reference model for $D_{r}$. The objective can be written as
@@ F.5 RMU
F.6 WHP Alternate Labels
In WHP (Eldan & Russinovich, 2023), the unlearning is performed through finetuning on alternate labels. The methods of producing alternate labels consists of two parts. Firstly, certain key terms (called “anchor terms”) are identified and replaced with generic terms according to a pre-set dictionary. For other tokens, this method hopes to identify the tokens that has the most positive change from a baseline model to a reinforced model that has been fitted to the unlearn data. Such increase are reversed according to the formula
@@ F.6 WHP Alternate Labels
Appendix G Relearn Text Itself is Uninformative
In this section, we show that our relearn set does not contain information or answers to the true evaluation target. Note that for TOFU, WHP, and the Synthetic dataset, we first finetune the base model so that it memorize the domain knowledge of these datasets. We perform the unlearning-relearning pipeline on top of that augmented model. To show the relearn set itself is insufficient to uncover the target knowledge, we directly finetune the base model (Phi-1.5 for TOFU, Llama-2-7b for WHP and Synthetic dataset) using the relearn set and test whether the resulting model provides us any useful information about the evaluation knowledge compared to the pretrained model. The result are shown in Table 9. On all three datasets, directly relearning with $D^{\prime}$ provides us no improvement on the evaluation task over the pretrained model. This suggests that knowledge within the evaluation set could not be retrieved directly from the relearn set. The effectiveness of relearning attack comes from the fact that unlearning fails to remove information within the forget set.
Dataset Evaluation metric Pretrained model Pretrained model finetuned with $D^{\prime}$
TOFU keyword search Accuracy 0% 0%
Synthetic keyword search Accuracy 0% 0%
WHP verbatim memorization Rouge-L score 0.1446 0.1407
Table 9: Comparison of model performance between the raw pretrained model and the pretrained model only finetuned with the relearn text.
Appendix H Algorithm
We provide the full algorithm of our attack in Algorithm 1.
Algorithm 1 Benign Relearning Attack
4:  for query $q\in S$ do
6:  end for
7:  return  Relearned model $w^{\prime}$ and score for every $q\in S$: $\{s_{q}\}$
Appendix I Detailed explanation of evaluation metrics
As target applications of unlearning may differ significantly (e.g., unlearning verbatim memorized text vs. unlearning general knowledge of a topic), it is common to consider different metrics/evaluation procedures for different unlearning tasks. In our work, we explore the following common evaluation metrics for specific unlearning tasks.
  1. 1.
    LLM-based evaluation. In applications such as removing hazardous or toxic knowledge, a reasonable goal for the adversary is to get high quality answers from the unlearned LLM using a harmful query. In this case, where quality is less well-defined through simple metrics, it is common to ask an existing LLM to provide a score. Following the recipe from Zheng et al. (2024), we use GPT-4 as the LLM Judge to provide a single score between 1-10 for every query-completion pair. A higher score means indicates that the completion more effectively answers the query.
  2. 2.
    Rouge-L. Another common application of unlearning is to mitigate verbatim memorization of copyrighted content. Under such settings, an adversary aims to see if the LLM will output the exact same copyrighted text used during training. Similar to Maini et al. (2024b), given the prefix $q$ and the true copyrighted continuation content $a$, we use Rouge-L score to measure the similarity between $a$ and the model output $o$. Specifically, we define $eval(o,q)=\text{Rouge}_{L}(a,o)$.
  3. 3.
    Keyword search. Finally, in certain applications the adversary may care about obtaining specific keywords such as names or ID numbers from the model. To evaluate whether such a goal is satisfied, one can simply check whether a keyword $k$ is contained in $o$. Specifically, $eval(o,q)$ could be defined as $eval(o,q)=\mathbf{1}_{k\in o}$.  [^2] Note that this is more flexible compared to exact string match based metric in prior work (Maini et al., 2024b) as we only care about the appearance of a few tokens.
Appendix J Additional presentation of results in Section 3
Figure 8: Line plot attack success rate for running different relearning steps on different unlearning checkpoints. Top: TOFU, Bottom: WHP.
We provide a line plot version of Figure 3 for an alternative presentation.